# LLaVA-CMoE: Towards Continual Mixture of Experts for Large Vision-Language Models

### v2

@@ 1 Introduction
Building on the aforementioned challenges in expert allocation and the mitigation of catastrophic forgetting, we propose an effective framework, LLaVA-CMoE, which requires no replay data while enabling truly continual learning. Our method facilitates adaptive parameter updates, enabling the network to autonomously determine when and where updates are necessary. These updates ensure that all previously acquired knowledge remains intact during the learning of new tasks. To tackle the challenge in expert allocation, we propose the Probe-Guided Knowledge Extension (PGKE) algorithm, which leverages a small amount of probe data from new tasks to monitor the behavior of experts newly added at each layer. Each expert functions as a probe; by observing the activation frequency of each expert probe, the network is able to independently determines whether additional parameters are required at each layer, avoiding unnecessary parameter growth in pu2023chatgpt . To address the challenge of catastrophic forgetting, we introduce the Probabilistic Task Locator (PTL) strategy. Rather than relying on a single shared router, which is susceptible to routing conflicts across tasks, we assign each task its own dedicated router. This design ensures that the knowledge and routing preferences for previous tasks can be preserved, as activating the corresponding router enables full recovery of prior performance. Since routers are lightweight compared to experts, this approach adds minimal computational overhead. However, during inference, the task identity is typically unknown and inputs are presented in a task-agnostic manner. To effectively identify the appropriate router for each input, we draw inspiration from pu2016variational ; pinheiro2021variational ; An2015VariationalAB and employ a reconstruction-based strategy. Specifically, we learn task-specific distributional representations under a VAE-based framework. During inference, the model evaluates the reconstruction error across all tasks and selects the one that best matches the input distribution. This enables PTL to autonomously infer the task category and dynamically route inputs to the most suitable experts, ensuring robust performance across diverse and unknown tasks.
We evaluate LLaVA-CMoE on the CoIN dataset chen2024coin , a benchmark tailored to assess continual learning performance across eight distinct VQA-based tasks. Through extensive quantitative and qualitative evaluations, along with comprehensive ablation studies, we demonstrate that our approach not only significantly mitigates catastrophic forgetting, but also promotes effective knowledge transfer and adaptation as new tasks are learned sequentially.
@@ 2 Related Work
Mixture of Experts. The Mixture of Experts (MoE) architecture riquelme2021scaling ; shen2023scaling ; mustafa2022multimodal employs specialized expert networks and a gating mechanism for efficient computation. Sparsely-gated MoE Shazeer_Mirhoseini_Maziarz_Davis_Le_Hinton_Dean_2017 ; lepikhin2020gshardscalinggiantmodels has shown strong performance in LLMs, such as Mixtral 8x7B jiang2024mixtralexperts , across diverse NLP tasks. Recent work liu2024moe ; yang2024moralmoeaugmentedlora ; chen2023octavius ; Luo2024MoELoRACL combines MoE with LoRA hu2022lora for more efficient training. MoE’s scalability has led to its adoption in continual learning, e.g., LEMoE wang2024lemoe adds experts at all layers for new tasks, and Lifelong-MoE chen2023lifelong trains new experts while freezing old ones. CoIN chen2024coin further explores MoELoRA, but existing methods still face parameter overhead and suboptimal robustness.
Large Language Models & PEFT. Recently, Large Language Models liu2023visual ; touvron2023llama ; chowdhery2023palm ; brown2020language have garnered widespread attention for their remarkable abilities in areas such as language generation, in-context learning, and reasoning. To enable data- and compute-efficient adaptation for specific downstream tasks, various Parameter-Efficient Fine-Tuning (PEFT) methods karimi2021compacter ; houlsby2019parameter ; li2021prefix have been introduced. Among these, LoRA hu2022lora stands out by representing weight updates through low-rank decomposition with two matrices, keeping the original weights frozen while training only the new update matrices. In this study, we combined LoRA with MoE for efficient continual MLLM fine-tuning.
@@ 3 Methodology
In this section, we first clarify the problem formulation of multimodal continual learning in Section 3.1. We then elaborate on the proposed LLaVA-CMoE and its core components: Probe-Guided Knowledge Expansion (PGKE) and Probabilistic Task Locator (PTL) in Section 3.2. Finally, we detail the training objectives in Section 3.3.
@@ 3.1 Problem Formulation
Let $\{\mathcal{T}_{1},\dots,\mathcal{T}_{N}\}$ be a set of $N$ tasks, where each task $\mathcal{T}_{i}$ has its training set $\mathbf{D}^{i}$ consisting of $n_{i}$ multimodal inputs $\mathbf{X}\in\{\mathbf{X}_{t}^{i},\mathbf{X}_{img}^{i},\mathbf{X}_{l}^{i}\}_{i=1}^{n_{i}}$. Here, $\mathbf{X}_{t}^{i}$, $\mathbf{X}_{img}^{i}$, and $\mathbf{X}_{l}^{i}$ represent the textual input (instruction), image, and the corresponding answer, respectively. In Continual Learning (CL), the model is trained sequentially on the $N$ tasks, and while training the $i$-th task $\mathcal{T}_{i}$, the model needs to maximize the probability $P_{i}$ through Next Token Prediction. Notably, while training the $i$-th task, it cannot access to the data of previous tasks $\{\mathcal{T}_{1},\dots,\mathcal{T}_{i-1}\}$. During the inference phase, we receive only task-agnostic data from either seen or unseen tasks. This practical constraint differentiates our approach from traditional task-incremental learning (Task-IL) paradigms, where explicit task identifiers are typically accessible during inference shi2024continual , thus significantly enhancing practical applicability in real-world scenarios where task boundaries may be ambiguous.
@@ 3.2.1 Mixture of Expert Layer
Our model is built upon a multimodal large language model (MLLM), such as LLaVA liu2023visual , in which the Mixture of Experts (MoE) is implemented by augmenting the Feed-Forward Network (FFN) modules within each Transformer block. Furthermore, each expert is constructed using a LoRA-style module, parameterized by $\{\mathbf{A}_{i},\mathbf{B}_{i}\}_{i=1}^{N_{e}}$, where $N_{e}$ denotes the number of experts. In the $h$-th block, given the multimodal token $\mathbf{X}_{h}$, the output token $\mathbf{X}_{h}^{out}$ is computed as follows:
$\mathbf{X}^{out}_{h}=\mathbf{W}_{0}\mathbf{X}_{h}+\sum_{i=1}^{N_{e}}\omega_{i}^{\prime}\mathbf{B}_{i}\mathbf{A}_{i}\mathbf{X}_{h},\quad\omega_{i}^{\prime}=\frac{\omega_{i}}{\sum^{N_{e}}_{j=1}{\omega_{j}}},\quad\omega_{i}=\exp(\mathbf{GX}_{h})\cdot\mathbb{I}[i\in\mathrm{topk}(\mathbf{GX}_{h})],$ (1)

### v3

@@ 1 Introduction
Building on the aforementioned challenges in expert allocation and the mitigation of catastrophic forgetting, we propose an effective framework, LLaVA-CMoE, which requires no replay data while enabling truly continual learning. Our method facilitates adaptive parameter updates, enabling the network to autonomously determine when and where updates are necessary. These updates ensure that all previously acquired knowledge remains intact during the learning of new tasks. To tackle the challenge in expert allocation, we propose the Probe-Guided Knowledge Extension algorithm, which leverages a small amount of probe data from new tasks to monitor the behavior of experts newly added at each layer. Each expert functions as a probe; by observing the activation frequency of each expert probe, the network is able to independently determines whether additional parameters are required at each layer, avoiding unnecessary parameter growth in pu2023chatgpt . To tackle catastrophic forgetting, we propose the Probabilistic Task Locator strategy. Instead of a single shared router prone to task conflicts, each task gets a dedicated router. This design preserves prior task knowledge and routing preferences—activating a router restores its performance. Lightweight routers introduce minimal overhead. During inference, task identity is unknown, and inputs are task-agnostic. Inspired by pu2016variational ; pinheiro2021variational ; An2015VariationalAB , we use a reconstruction-based strategy: via a VAE framework, we learn task-specific distributions. The model evaluates reconstruction errors to find the best-matching task distribution, allowing PTL to infer tasks, route inputs dynamically, and maintain robust performance on unknown tasks. We evaluate LLaVA-CMoE on the CoIN dataset chen2024coin , a benchmark tailored to assess continual learning performance across eight distinct VQA-based tasks. Through extensive quantitative and qualitative evaluations, along with comprehensive ablation studies, we demonstrate that our approach not only significantly mitigates catastrophic forgetting, but also promotes effective knowledge transfer and adaptation as new tasks are learned sequentially.
@@ 2 Related Work
Mixture of Experts. The MoE architecture riquelme2021scaling ; shen2023scaling ; mustafa2022multimodal employs specialized expert networks and a gating mechanism for efficient computation. Sparsely-gated MoE Shazeer_Mirhoseini_Maziarz_Davis_Le_Hinton_Dean_2017 ; lepikhin2020gshardscalinggiantmodels has shown strong performance in LLMs, such as Mixtral 8x7B jiang2024mixtralexperts , across diverse NLP tasks. Recent work liu2024moe ; yang2024moralmoeaugmentedlora ; chen2023octavius ; Luo2024MoELoRACL combines MoE with LoRA hu2022lora for more efficient training. MoE’s scalability has led to its adoption in continual learning, e.g., LEMoE wang2024lemoe adds experts at all layers for new tasks, and Lifelong-MoE chen2023lifelong trains new experts while freezing old ones. CoIN chen2024coin further explores MoELoRA, but existing methods still face parameter overhead and suboptimal robustness.
Large Language Models & PEFT. Recently, Large Language Models liu2023visual ; touvron2023llama ; chowdhery2023palm ; brown2020language have garnered widespread attention for their remarkable abilities in areas such as language generation, in-context learning, and reasoning. To enable data- and compute-efficient adaptation for specific downstream tasks, various PEFT methods karimi2021compacter ; houlsby2019parameter ; li2021prefix have been introduced. Among these, LoRA hu2022lora stands out by representing weight updates through low-rank decomposition with two matrices, keeping the original weights frozen while training only the new update matrices. In this study, we combined LoRA with MoE for efficient continual MLLM fine-tuning.
@@ 3 Methodology
In this section, we first clarify the problem formulation of multimodal continual learning in Section 3.1. We then elaborate on the proposed LLaVA-CMoE and its core components: Probe-Guided Knowledge Expansion and Probabilistic Task Locator in Section 3.2. Finally, we detail the training objectives in Section 3.3.
@@ 3.1 Problem Formulation
Let $\{\mathcal{T}_{1},\dots,\mathcal{T}_{N}\}$ be a set of $N$ tasks, where each task $\mathcal{T}_{i}$ has its training set $\mathbf{D}^{i}$ consisting of $n_{i}$ multimodal inputs $\mathbf{X}\in\{\mathbf{X}_{t}^{i},\mathbf{X}_{img}^{i},\mathbf{X}_{l}^{i}\}_{i=1}^{n_{i}}$. Here, $\mathbf{X}_{t}^{i}$, $\mathbf{X}_{img}^{i}$, and $\mathbf{X}_{l}^{i}$ represent the textual input (instruction), image, and the corresponding answer, respectively. In Continual Learning, the model is trained sequentially on the $N$ tasks, and while training the $i$-th task $\mathcal{T}_{i}$, the model needs to maximize the probability $P_{i}$ through Next Token Prediction. Notably, while training the $i$-th task, it cannot access to the data of previous tasks $\{\mathcal{T}_{1},\dots,\mathcal{T}_{i-1}\}$. During the inference phase, we receive only task-agnostic data from either seen or unseen tasks. This practical constraint differentiates our approach from traditional task-incremental learning (Task-IL) paradigms, where explicit task identifiers are typically accessible during inference shi2024continual , thus significantly enhancing practical applicability in real-world scenarios where task boundaries may be ambiguous.
@@ 3.2.1 Mixture of Expert Layer
Our model is built upon a multimodal large language model (MLLM), such as LLaVA liu2023visual , in which the MoE is implemented by augmenting the Feed-Forward Network (FFN) modules within each Transformer block. Furthermore, each expert is constructed using a LoRA module, parameterized by $\{\mathbf{A}_{i},\mathbf{B}_{i}\}_{i=1}^{N_{e}}$, where $N_{e}$ denotes the number of experts. In the $h$-th block, given the multimodal token $\mathbf{X}^{h}$, the output token $\mathbf{X}_{out}^{h}$ is computed as follows:
$\mathbf{X}_{out}^{h}=\mathbf{W}_{0}\mathbf{X}^{h}+\sum_{i=1}^{N_{e}}\omega_{i}^{\prime}\mathbf{B}_{i}\mathbf{A}_{i}\mathbf{X}^{h},\quad\omega_{i}^{\prime}=\frac{\omega_{i}}{\sum^{N_{e}}_{j=1}{\omega_{j}}},\quad\omega_{i}=\exp(\mathbf{GX}^{h})\cdot\mathbb{I}[i\in\mathrm{topk}(\mathbf{GX}^{h})],$ (1)